# Concepts in Motion: Temporal Bottlenecks for Interpretable Video Classification

### v1

@@ Abstract
Conceptual models such as Concept Bottleneck Models (CBMs) have driven substantial progress in improving interpretability for image classification by leveraging human-interpretable concepts. However, extending these models from static images to sequences of images, such as video data, introduces a significant challenge due to the temporal dependencies inherent in videos, which are essential for capturing actions and events. In this work, we introduce MoTIF (Moving Temporal Interpretable Framework), an architectural design inspired by a transformer that adapts the concept bottleneck framework for video classification and handles sequences of arbitrary length. Within the video domain, concepts refer to semantic entities such as objects, attributes, or higher-level components (e.g., “bow,” “mount,” “shoot”) that reoccur across time—forming motifs collectively describing and explaining actions. Our design explicitly enables three complementary perspectives: global concept importance across the entire video, local concept relevance within specific windows, and temporal dependencies of a concept over time. Our results demonstrate that the concept-based modeling paradigm can be effectively transferred to video data, enabling a better understanding of concept contributions in temporal contexts while maintaining a competitive performance. Code available at github.com/patrick-knab/MoTIF.
@@
1 Introduction
Modern deep learning models already achieve outstanding results in video understanding tasks such as video classification, action recognition, and event detection (Liu et al., 2021; Bertasius et al., 2021). Despite their success, these models are commonly perceived as black boxes since their internal workings are not interpretable in a way that reveals their decision-making process (Molnar et al., 2020; Knab et al., 2025b). Concept Bottleneck Models (CBMs) (Koh et al., 2020) address this issue by enforcing an intermediate bottleneck layer of human-understandable concepts, which are then used by a linear classifier to generate the final prediction.
While CBMs have been extensively studied in the image domain (Prasse et al., 2025; Yang et al., 2023; Sun et al., 2025; Schrodi et al., 2025), their extension to video remains largely unexplored (Jeyakumar et al., 2022). Videos differ from images in that they contain a temporal component: concepts evolve over time and many actions cannot be deduced from a single frame (Lee et al., 2025; Chen et al., 2025). In addition, the length of a video is not fixed, making the adaptation of existing CBM architectures non trivial, since they always require the same embedding dimensionality. Transformer-based models (Vaswani et al., 2017) are powerful for modeling such long-range dependencies (Bertasius et al., 2021), but their dense feature mixing hinders clear attribution at the concept level, making their interpretation infeasible (Hao et al., 2021; Molnar et al., 2020).
In this work, we introduce MoTIF (Moving Temporal Interpretable Framework), a concept bottleneck model tailored for video classification. MoTIF builds on transformer-inspired blocks and introduces a per-channel temporal self-attention (diagonal attention) mechanism that isolates temporal reasoning for each concept via depthwise $1{\times}1$ convolutions. To illustrate how MoTIF extends beyond static images, Fig. 1 shows our framework: it processes a video, tracks its concepts through time, and explains which concepts drive the final prediction. In this example, the model identifies reoccurring semantic motifs such as a bow being mounted and the arrow being shot by Legolas (Lord of the Rings), which together form the basis for recognizing the higher-level action “shooting a bow.” Temporal dependency maps further reveal that specific concepts primarily attend to characteristic frames, i.e., the moments when the bow is mounted and the arrow released. As this example shows, the design preserves concept separation across time and enables analysis of both global concept contributions over an entire video and local relevance within specific frames.
@@ 1 Introduction
  • A CBM framework, MoTIF, for video sequences that supports arbitrary-length inputs and integrates seamlessly with vision–language backbones.
  • Per-channel temporal self-attention that preserves concept independence within transformer blocks and models temporal dynamics on a per-concept basis.
  • MoTIF is the first method to enable three complementary explanation modes: (i) global concept relevance via log-sum-exp (LSE) pooling, (ii) localized temporal explanations using windowed concept attributions, and (iii) attention-based temporal maps that visualize how a concept channel distributes its focus across time.
Extensive experiments and ablations (window size, backbone selection, temperature $\tau$, full vs. diagonal attention, etc.) demonstrate that MoTIF achieves robust performance while providing fine-grained and faithful explanations.
@@ 2 Method
In this section, we present the MoTIF framework for interpretable video classification. The goal is to preserve concept-level transparency while leveraging temporal modeling capabilities.
@@
MoTIF is a CBM with a transformer architecture for video classification that operates directly on sequences of per-window concept activations (Figure 2). A video is represented by $T$ temporal windows, each described by $C$ scalar concept activations, with $k=|C|$; the input is $X\in\mathbb{R}^{T\times C}$.
@@
The framework enforces strict per-concept interpretability: temporal dependencies are modeled independently for each concept channel, so that recurrent temporal patterns—motifs—remain attributable to individual concepts. After temporal processing, per-concept activations are refined via a nonnegative affine transformation, and a classifier produces per-time-step logits. Video-level predictions are obtained via log-sum-exp pooling, which also yields a time-importance profile for additional explanation. In contrast, a conventional transformer mixes channels during attention (Vaswani et al., 2017). For comparison, we also evaluate a variant with full multi-head attention, but emphasize that this removes explicit concept attribution. A detailed analysis of this and other design choices is provided in Section 3.2.2 and Appendix C. For completeness, an algorithmic overview of the framework and its subparts is included in Appendix E.
Channels as concepts.
@@
We use the term channel to denote one dimension of the concept space. Each video $n$ yields $y^{(n)}\in\mathbb{R}$ classes, $X^{(n)}\in\mathbb{R}^{T_{n}\times C}$ concept activations, where the second axis ($C$) indexes concepts; each slice $X^{(n)}_{:,c}\in\mathbb{R}^{T_{n}}$ corresponds to the temporal activation sequence of concept $c\in C$. Unlike CNN channels, these dimensions are semantically interpretable by design.
Notation.
Capital and lowercase letters denote matrices and vectors, respectively. During training, a batch of videos is represented as
$X\in\mathbb{R}^{B\times T\times C},$ (1)
@@
where $B$ is the batch size, $T$ is the (padded) sequence length, and $C$ is the number of concepts. Thus $X_{:,:,c}\in\mathbb{R}^{B\times T}$ contains the activations of concept $c$ across the batch.
@@ 2.1 Transformer Bottleneck Model
Video and concept embeddings. We employ an image–text aligned backbone (Radford et al., 2021) to map frames into a shared embedding space. Window embeddings are obtained by a random frame in the window or via a video-adapted CLIP variant (see Appendix A.3), and concept activations are computed as cosine similarities to a defined concept bank, producing $\{X^{(n)}\}_{n=1}^{N}$ with $X^{(n)}\in\mathbb{R}^{T_{n}\times C}$. The concept bank $C$ is constructed from natural language descriptions of human-interpretable actions and objects. To generate a diverse and semantically rich vocabulary, we use a large language model (LLM) (OpenAI et al., 2024) (GPT-5) to propose candidate concepts, following Yang et al. (2023) who leverage LLMs for automated concept set generation. Since concept generation is not the main focus of this work, we adopt this approach without further refinement.
@@ 2.1.1 Per-Channel Temporal Self-Attention
$Q,K,V\in\mathbb{R}^{T\times C},Q_{:,c}=x_{:,c}\,\theta_{Q}^{(c)},\;K_{:,c}=x_{:,c}\,\theta_{K}^{(c)},\;V_{:,c}=x_{:,c}\,\theta_{V}^{(c)},$ (2)
with $\theta_{Q,K,V}^{(c)}\in\mathbb{R}$. Here $\theta_{Q}^{(c)},\theta_{K}^{(c)},\theta_{V}^{(c)}$ are channel-specific scalars applied uniformly across all $T$, not temporal filters. Equivalently, each projection uses a depthwise kernel $\Theta_{Q},\Theta_{K},\Theta_{V}\in\mathbb{R}^{C\times 1\times 1}$, where $\theta_{*}^{(c)}$ is the $c$-th depthwise filter. Thus, temporal information is preserved and channels remain independent, without cross-channel mixing. Attention scores are computed per concept as $W_{c,t,u}\;=\;Q_{t,c}\,K_{u,c}$, where $t$ denotes the query time step and $u$ the key/value time step. A softmax over $u$ yields attention weights $W\in\mathbb{R}^{C\times T\times T}$, so that each concept decides which of its past or future activations to attend to. The output is obtained as the weighted sum over $V$:
@@
The difference between full and diagonal attention is visualized in Equation 4 as the structure of the concept–concept attention matrix. That is the reason why we call this mechanism diagonal attention: each channel learns to construct its own temporal filter by weighting past activations differently at each step. Unlike fixed convolutional kernels, the attention weights adapt to the input sequence, yet the restriction to depthwise (per-concept) projections guarantees that evidence for one concept does not leak into another.
$\begin{array}[]{cc}\underbrace{\begin{bmatrix}(c_{1}\!\to\!c_{1})&(c_{1}\!\to\!c_{2})&\cdots&(c_{1}\!\to\!c_{C})\\ (c_{2}\!\to\!c_{1})&(c_{2}\!\to\!c_{2})&\cdots&(c_{2}\!\to\!c_{C})\\ \vdots&\vdots&\ddots&\vdots\\ (c_{C}\!\to\!c_{1})&(c_{C}\!\to\!c_{2})&\cdots&(c_{C}\!\to\!c_{C})\end{bmatrix}}_{\text{Full attention}}&\underbrace{\begin{bmatrix}(c_{1}\!\to\!c_{1})&&&\\ &(c_{2}\!\to\!c_{2})&&\\ &&\ddots&\\ &&&(c_{C}\!\to\!c_{C})\end{bmatrix}}_{\text{Diagonal attention}}\end{array}$ (4)
The block concludes with per-channel normalization and a lightweight feed-forward network (two depthwise $1{\times}1$ convolutions with GELU and dropout). For MoTIF, stacking more blocks does not add richer hierarchical abstractions, because cross-concept interactions (which normally grow expressivity in deep transformers) are deliberately suppressed.
@@ 2.1.3 Classification Head
Training objective.
The model is trained with class-weighted cross-entropy on $\hat{\ell}$, complemented with two regularizers: an $\ell_{1}$ penalty on $W$ to encourage sparsity, and an activation sparsity penalty on $Z$:
$\mathcal{L}\;=\;\mathrm{CE}(\hat{\ell},y)+\lambda_{\ell_{1}}\,\lVert W_{k}\rVert_{1}+\lambda_{\mathrm{sparse}}\;\frac{1}{(\sum_{t}m_{t})C}\sum_{t,c}m_{t}|Z_{t,c}|.$ (6)

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) enable interpretable image classification by structuring predictions around human-understandable concepts, but extending this paradigm to video remains challenging due to the difficulty of extracting concepts and modeling them over time. In this paper, we introduce MoTIF (Moving Temporal Interpretable Framework), a transformer-based concept architecture that operates on sequences of temporally grounded concept activations, by employing per-concept temporal self-attention to model when individual concepts recur and how their temporal patterns contribute to predictions. Central to the framework is an agentic concept discovery module to automatically extract object- and action-centric textual concepts from videos, yielding temporally expressive concept sets without manual supervision. Across multiple video benchmarks, this combination substantially narrows the performance gap between interpretable and black-box video models while maintaining faithful and temporally grounded concept explanations. Code available at github.com/patrick-knab/MoTIF.
Machine Learning, ICML
1 Introduction
Modern deep learning models achieve outstanding results in video tasks such as video classification, action recognition, and event detection (Liu et al., 2021; Bertasius et al., 2021). Despite their success, these models are commonly perceived as black boxes since their internal workings are not interpretable (Molnar et al., 2020; Knab et al., 2025b) and this opacity can limit deployment in medical domains, where interpretability may matter more than peak accuracy. Concept Bottleneck Models (CBMs) (Koh et al., 2020) address this issue by enforcing an intermediate bottleneck layer of human-understandable concepts, which are then used by a linear classifier to generate the final prediction.
Figure 1: Overview of MoTIF. Video classification pipeline with agentic concept discovery and its three explanation modes.
@@ 1 Introduction
While CBMs have been extensively studied in the image domain (Prasse et al., 2025; Yang et al., 2023; Sun et al., 2025; Schrodi et al., 2025), their extension to video remains largely unexplored (Jeyakumar et al., 2022). Videos differ from images in that they contain a temporal component: concepts evolve over time, and many actions cannot be inferred from a single frame (Lee et al., 2025b; Chen et al., 2025). Importantly, the challenge of variable video length is orthogonal to modeling long-range dependencies. While transformers (Vaswani et al., 2017) excel at capturing such dependencies (Bertasius et al., 2021), their dense temporal feature mixing obscures concept-level attributions, limiting their suitability for interpretable concept bottleneck modeling (Hao et al., 2021; Molnar et al., 2020). This highlights the need for architectures that handle variable-length inputs while preserving concept-level interpretability over time.
In this work, we introduce MoTIF (Moving Temporal Interpretable Framework), a concept bottleneck model tailored for video classification. The CBM is not designed to replace high-capacity black-box video models, but to make temporal concept reasoning explicit and inspectable. MoTIF builds on transformer-inspired blocks and introduces a per-channel temporal self-attention (diagonal attention) mechanism that isolates temporal reasoning for each concept to enable interpretation. In addition, we utilize an agentic concept discovery approach to generate object- and action-based textual, domain-specific concepts (Fig. 1). This setup removes the need for manually curated concept vocabularies, which is particularly valuable in video domains where relevant objects and actions vary substantially across datasets. To illustrate how MoTIF extends beyond static images, Fig. 2 shows our framework: it processes a video, tracks its concepts through time, and explains which concepts drive the final prediction. In this example, the model identifies recurring semantic motifs such as a bow being mounted and the arrow being shot by Legolas (Lord of the Rings), which together form the basis for recognizing the higher-level action “shooting a bow.” Temporal dependency maps further reveal that specific concepts primarily attend to characteristic frames, i.e., the moments when the bow is mounted and the arrow released. As this example shows, the design preserves concept separation across time and enables analysis of both global concept contributions over an entire video and local relevance within specific frames.
@@
  • MoTIF, a Video CBM with per-channel temporal self-attention that preserves concept independence within Transformer blocks while modeling temporal dynamics at the level of individual concepts.
  • An unsupervised agentic concept discovery pipeline that extracts object and action concepts directly from video and substantially improves the performance of interpretable video CBMs, narrowing the gap to black-box models.
  • MoTIF enables three complementary explanation modes: (i) global concept relevance via log-sum-exp (LSE) pooling, (ii) localized temporal explanations using windowed concept attributions, and (iii) attention-based temporal maps over concept activations.
@@ 2 Method
In this section, we present the agentic concept discovery and the MoTIF framework for interpretable video classification. The goal is to preserve concept-level transparency while leveraging temporal modeling capabilities.
@@
MoTIF is a CBM with a transformer architecture for video classification that operates directly on sequences of per-window concept activations (Figure 2). A video is represented as a sequence of $T$ temporal windows, each encoded by $C$ scalar concept activations (one per concept), where $C=|\mathcal{C}|$ is given by the concept bank produced by the agentic discovery module.
@@
The framework enforces strict and intentional per-concept interpretability: temporal dependencies are modeled independently for each concept channel, recurrent temporal patterns—motifs—remain attributable to individual concepts. After temporal processing, per-concept activations are refined via a nonnegative affine transformation, and a classifier produces per-time-step logits. Video-level predictions are obtained via log-sum-exp pooling, which also yields a time-importance profile for additional explanation. In contrast, a conventional transformer mixes channels during attention (Vaswani et al., 2017)  [^1] We also evaluate full multi-head attention, but emphasize that this removes explicit concept attribution. A detailed analysis and other design choices are provided in Section 3.2.2 and Appendix C. For completeness, Appendix D provides an algorithmic overview of MoTIF, including its main components and the concept discovery module.
@@
Channels as concepts. We use the term channel to denote one dimension of the concept space. Each video $n$ yields $y^{(n)}\in\mathbb{R}$ classes, $X^{(n)}\in\mathbb{R}^{T_{n}\times C}$ concept activations, where the second axis ($C$) indexes concepts; each slice $X^{(n)}_{:,c}\in\mathbb{R}^{T_{n}}$ corresponds to the temporal activation sequence of concept $c\in C$. Unlike CNN channels, these dimensions are semantically interpretable by design.
@@
Notation. Capital and lowercase letters denote matrices and vectors, respectively. During training, a batch of videos is represented as $X\in\mathbb{R}^{B\times T\times C}$, where $B$ is the batch size, $T$ is the (padded) sequence length, and $C$ is the number of concepts. Thus $X_{:,:,c}\in\mathbb{R}^{B\times T}$ contains the activations of concept $c$ across the batch.
2.1 Agentic Concept Discovery
The agentic concept discovery stage converts raw visual inputs into structured, interpretable concept activations that serve as input to the MoTIF backbone. While formulated for video, images are naturally handled as the special case of a single temporal window.
Temporal windowing. Given a video $X=\{I_{t}\}_{t=1}^{T}$, we partition it into $n$ temporal windows
$\mathcal{W}(x)=\{w_{k}\}_{k=1}^{n},\qquad w_{k}=\{I_{t}\}_{t=s_{k}}^{s_{k}+W-1},$ (1)
where $W$ denotes the window length and $\{s_{k}\}$ are the start indices. The number of windows $K$ is determined by the video duration, window size, and overlap. Smaller windows favor short-lived, fine-grained events, while larger windows emphasize persistent semantic content. This design provides explicit control over the temporal resolution at which concepts are discovered.
Agentic concept proposal. Each window $w_{k}$ is processed independently by a vision–language agent $A_{\mathrm{vlm}}$, which generates natural-language concept candidates
$(\mathcal{C}^{\text{obj}}_{w_{k}},\mathcal{C}^{\text{act}}_{w_{k}})=A_{\mathrm{vlm}}(w_{k}),$ (2)
capturing objects, attributes, actions, and interactions. This step is task-agnostic and does not rely on any predefined concept vocabulary (Appendix E.1).
Concept and video embeddings. The global concept bank $\mathcal{C}$ is constructed from the union of all discovered textual object and action concepts. To ensure semantic diversity and remove near-duplicates, concepts are embedded using the same image–text aligned backbone as MoTIF (CLIP), and filtered by cosine similarity: concepts with pairwise similarity greater than $0.9$ are removed. For each window, depending on the backbone, visual embeddings are obtained either from a representative frame or a video-adapted CLIP variant (Appendix A.3). Concept activations are computed as cosine similarities between window embeddings and concept embeddings, yielding per-video activation matrices
$X^{(n)}\in\mathbb{R}^{T_{n}\times C},$ (3)
where each channel corresponds to one interpretable concept and is processed independently by MoTIF.
@@ 2.2.1 Per-Channel Temporal Self-Attention
$Q,K,V$ $\in\mathbb{R}^{T\times C},Q_{:,c}$ $=x_{:,c}\,\theta_{Q}^{(c)},$ (4)
$K_{:,c}$ $=x_{:,c}\,\theta_{K}^{(c)},V_{:,c}$ $=x_{:,c}\,\theta_{V}^{(c)}.$
Here $\theta_{Q}^{(c)},\theta_{K}^{(c)},\theta_{V}^{(c)}\in\mathbb{R}$ are channel-specific scalars applied uniformly across all $T$, not temporal filters. Equivalently, each projection uses a depthwise kernel $\Theta_{Q},\Theta_{K},\Theta_{V}\in\mathbb{R}^{C\times 1\times 1}$, where $\theta_{*}^{(c)}$ is the $c$-th depthwise filter. Thus, temporal information is preserved and channels remain independent, without cross-channel mixing.
Attention scores are computed per concept as $W_{c,t,u}\;=\;Q_{t,c}\,K_{u,c}$, where $t$ denotes the query time step and $u$ the key/value time step. A softmax over $u$ yields attention weights $W\in\mathbb{R}^{C\times T\times T}$, so that each concept decides which of its past or future activations to attend to. The output is obtained as the weighted sum over $V$:
@@
The difference between full and diagonal attention is visualized in Equation 6 as the structure of the concept–concept attention matrix. This is why we call this mechanism diagonal attention: each channel learns to construct its own temporal filter by weighting past activations differently at each step. Unlike fixed convolutional kernels, the attention weights adapt to the input sequence, yet the restriction to depthwise (per-concept) projections guarantees that evidence for one concept does not leak into another.
$\begin{array}[]{cc}\underbrace{\left[\begin{smallmatrix}(c_{1}\to\ c_{1})&\cdots&(c_{1}\to\ c_{C})\\ \vdots&\ddots&\vdots\\ (c_{C}\to\ c_{1})&\cdots&(c_{C}\to\ c_{C})\end{smallmatrix}\right]}_{\text{Full attention}}&\underbrace{\left[\begin{smallmatrix}(c_{1}\to\ c_{1})&&\\ &\ddots&\\ &&(c_{C}\to\ c_{C})\end{smallmatrix}\right]}_{\text{Diagonal attention}}\end{array}$ (6)
The block concludes with per-channel normalization and a lightweight feed-forward network (two depthwise $1{\times}1$ convolutions with GELU and dropout). In this work, we deliberately isolate temporal reasoning per concept to preserve attribution. Importantly, this is a design choice rather than a limitation of the framework: as shown in Section 3.2.2 enabling cross-concept interactions improves accuracy at the cost of interpretability, highlighting a controllable trade-off rather than a hard restriction.
Architectural extension. For most experiments in this paper, we report the results of MoTIF using diagonal attention within a standard transformer architecture. However, as shown by (Bertasius et al., 2021), separating spatial and temporal attention can further enhance performance. Therefore, we additionally evaluate MoTIF when extended to a space-time transformer architecture (MoTIF-ST), as detailed in Appendix A.4, demonstrating that MoTIF is not restricted to a single transformer design.
@@ 2.2.3 Classification Head
Training objective. The model is trained with class-weighted cross-entropy on $\hat{\ell}$, complemented with two regularizers: an $\ell_{1}$ penalty on $W$ to encourage sparsity, and an activation sparsity penalty on $Z$:
$\mathcal{L}$ $=\mathrm{CE}(\hat{\ell},y)+\lambda_{\ell_{1}}\,\lVert W_{k}\rVert_{1}$ (8)
$\quad+\lambda_{\mathrm{sparse}}\;\frac{1}{(\sum_{t}m_{t})C}\sum_{t,c}m_{t}\lvert Z_{t,c}\rvert$.
@@ 2.3 Explanation
Table 1: Performance comparison (% Top-1 accuracy). Mean $\pm$ standard deviation on train-test splits across datasets, using agentic concept discovery for MoTIF, and comparing against zero-shot, global CBM, and black-box baselines.
Method Breakfast HMDB51 UCF101 SSv2
Zero-shot CLIP-RN/50 (Radford et al., 2021) 18.6 $\pm$ 2.6 29.8 $\pm$ 0.5 57.2 $\pm$ 0.9 0.8
CLIP-ViT-L/14 (Radford et al., 2021) 31.1 $\pm$ 4.7 45.7 $\pm$ 0.1 70.6 $\pm$ 0.5 0.9
PE-L/14 (Bolya et al., 2025) 41.4 $\pm$ 7.0 56.7 $\pm$ 0.6 74.6 $\pm$ 0.9 2.2
PE-G/14 (Bolya et al., 2025) 47.4 $\pm$ 5.4 60.7 $\pm$ 1.0 74.6 $\pm$ 0.9 2.2
Global CBM CLIP-RN/50 (Radford et al., 2021) 36.9 $\pm$ 7.7 61.8 $\pm$ 1.7 84.1 $\pm$ 0.8 17.0
CLIP-ViT-B/32 (Radford et al., 2021) 38.3 $\pm$ 9.5 62.6 $\pm$ 0.8 86.4 $\pm$ 0.6 18.2
CLIP-ViT-L/14 (Radford et al., 2021) 57.0 $\pm$ 8.0 71.0 $\pm$ 1.1 93.4 $\pm$ 0.7 22.0
PE-L/14 (Bolya et al., 2025) 72.4 $\pm$ 8.3 76.4 $\pm$ 0.8 96.3 $\pm$ 0.1 31.3
PE-G/14 (Bolya et al., 2025) 75.8 $\pm$ 7.1 77.8 $\pm$ 0.8 97.5 $\pm$ 0.4 33.6
MoTIF (ours) MoTIF (RN/50) 55.1 $\pm$ 7.1 66.2 $\pm$ 0.5 86.7 $\pm$ 0.6 20.0
MoTIF (ViT-B/32) 52.7 $\pm$ 5.8 68.5 $\pm$ 1.0 88.5 $\pm$ 0.6 20.7
MoTIF (ViT-L/14) 71.0 $\pm$ 6.2 76.1 $\pm$ 0.5 94.8 $\pm$ 0.5 25.8
MoTIF-ST (ViT-L/14) 72.6 $\pm$ 6.5 75.8 $\pm$ 0.6 94.8 $\pm$ 0.4 27.7
MoTIF (PE-L/14) 83.2 $\pm$ 6.2 81.8 $\pm$ 0.6 97.0 $\pm$ 0.3 37.3
MoTIF-ST (PE-L/14) 85.4 $\pm$ 6.3 80.8 $\pm$ 1.0 97.2 $\pm$ 0.2 39.6
MoTIF (PE-G/14) 87.5 $\pm$ 4.9 83.0 $\pm$ 0.6 98.0 $\pm$ 0.2 40.4
MoTIF-ST (PE-G/14) 87.3 $\pm$ 7.1 82.1 $\pm$ 1.0 98.4 $\pm$ 0.3 41.9
Black- Box TSM (Lin et al., 2019) 59.1 73.5 95.9 61.7
No frame left behind (Liu et al., 2021) 62.0 73.4 96.4 62.7
VideoMAE V2 (Wang et al., 2023) 88.1 99.6 76.8